# Leveraging Multi-facet Paths for Heterogeneous Graph Representation Learning

### v2

@@ 5.5. Visualization
To evaluate embedding quality, we visualize node representations from various HGNNs on the DBLP dataset using t-SNE (Van der Maaten and Hinton, 2008). Figure 6 shows t-SNE plots for HAN (Wang et al., 2019), MAGNN (Fu et al., 2020), HeCo (Wang et al., 2021), HMSG (Guan et al., 2023), MECCH (Fu and King, 2024), and MF2Vec, with colors representing author classes. MF2Vec demonstrates superior clustering within classes and clear decision boundaries, outperforming the more scattered intra-class distributions and weaker inter-class separations of other HGNNs.
6. Study of MF2Vec
To analyze the impact of intrinsic characteristics of MF2Vec with respect to various features including train rati0, the number facets, we additionally conducted extensive experiments.
Dataset | Yelp | Movielens | DBLP | ACM | IMDB | Freebase | ||||||||||||
Random. | 0.8670.018 | 0.7970.003 | 0.8080.015 | 0.8710.006 | 0.7950.006 | 0.8030.005 | 0.9890.003 | 0.9350.009 | 0.9310.01 | 0.9490.009 | 0.8240.013 | 0.8130.012 | 0.7510.005 | 0.5880.004 | 0.5710.003 | 0.8250.020 | 0.6960.017 | 0.6280.042 |
6.1. Train Ratio ()
@@
As shown in Figure 7, MF2Vec consistently achieves the highest performance across varying training set ratios, demonstrating its robustness and adaptability. While the performance of all models declines as the training ratio decreases, MF2Vec remains its superiority in most cases and retains competitive in the rest. At the lowest ratio of 0.2, it occasionally shows comparable performance to other models. HAN and one baseline were excluded due to significantly lower performance, ensuring clearer visualization. These results highlight MF2Vec’s effectiveness in handling sparse training data.
6.2. Number of Facets ()
@@
6.3. Role of Facet Selection
As MF2Vec selects the optimal facet for each path in a learnable manner, we conduct an additional experiment to validate its facet selection strategy. Table 8 compares different methods for selecting facet information as edge weights. Our GCN-based results show that the proposed method, using Gumbel-Softmax and attention mechanisms, outperforms all alternatives across datasets. Random facet selection reduces performance, and omitting facets entirely leads to the lowest results. These findings highlight the importance of facet-related information for effective graph learning and the proposed method’s ability to capture complex structures and hidden relationships.
7. Case Study
To evaluate facet balance among node types, we analyzed their distribution as shown in Figure 9. The heatmaps reveal equitable attention weights across node types, with each column representing a specific node and highlighting diverse facet distributions. For example, the Paper node “P1612” emphasizes Facet 1, while the Author node “A3958” distributes attention evenly across facets, reflecting varying significance assigned to facets based on connections. This analysis confirms that nodes either focus on specific facets or distribute attention uniformly, with no strong bias toward any facet, validating the importance of modeling shared facets in heterogeneous graphs.
@@ 8. Conclusion
In this study, we introduce MF2Vec, a novel heterogeneous graph neural network model, which generates node representations using multi-facet vectors from various paths, regardless of node types. MF2Vec captures complex latent semantics through these multi-faceted paths, offering a node-type-agnostic embedding approach. Evaluations across six datasets demonstrate that MF2Vec outperforms leading baseline models in both performance and time complexity. Future work could focus on automating path selection within heterogeneous graph neural networks using learnable methods to improve upon current random pre-selection methods, highlighting the potential of dynamic multi-facet path utilization in graph networks.
Appendix
A. Baselines Code and Datasets
MF2Vec is implemented using PyTorch. Table 9 lists the URLs of the authors’ implementations of the compared methods, and Table 10 provides the URLs where the datasets can be downloaded.
B. Baselines
We compare MF2Vec against state-of-the-art relation-based HGNNs and metapath-based HGNNs , categorized as follows:
Relation-Based HGNNs.
RGCN: Introduces a relational graph convolutional network that assigns weights to each relation type, enabling the learning of structural information in heterogeneous graphs.
RGAT: Applies attention mechanisms to each relation type in relational graphs, effectively integrating information in heterogeneous graphs.
HGT: Utilizes a modified transformer architecture that considers various node and edge types in heterogeneous graphs to learn complex relationships.
HetSANN: Applies attention mechanisms to learn structural information between nodes in heterogeneous graphs without relying on metapaths.
Simple-HGN: Provides efficient and effective node embeddings in heterogeneous graphs through a simplified structure.
Metapath-Based HGNNs.
MP2Vec: Proposes a node embedding technique in heterogeneous networks using metapaths to effectively capture diverse relationships.
HAN: Develops a hierarchical attention mechanism-based model to learn the importance of nodes in heterogeneous graphs using metapaths.
MAGNN: Enhances node embedding performance in heterogeneous graphs by leveraging intermediate node information within metapaths.
HeCo: Combines self-supervised learning and contrastive learning in heterogeneous graphs to improve node representation learning performance.
MHNF: A framework that automatically extracts metapaths in heterogeneous networks and utilizes them for node embedding learning.
HGMAE: Utilizes masked autoencoders to learn robust node representations in heterogeneous graphs.
HMSG: Integrates multi-scale structural information in heterogeneous graphs to enhance node representation learning performance.
MECCH: Improves node embedding performance in heterogeneous graphs by leveraging contextual information of metapaths.
C. Stability of Performance Across Metapaths
Tables 11, 12, and 13 present performance metrics for the top three models and MF2Vec across the DBLP, ACM, and IMDB datasets, respectively, based on a single metapath. Across all datasets, MF2Vec consistently outperforms other models, achieving the highest Mi-F1, Ma-F1, and AUC scores for all metapaths. On the DBLP dataset (Table 11), MF2Vec excels across all metrics. Similarly, on the ACM dataset (Table 12), it achieves the best performance for both PAP and PSP metapaths. On the IMDB dataset (Table 13), MF2Vec again demonstrates its robustness and consistency, achieving the highest scores across all metapaths.
DBLP | ||||||||||||
Metapath | APTPA | APCPA | APA | ALL | ||||||||
Metric | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC |
MAGNN | 0.670 | 0.685 | 0.912 | 0.918 | 0.925 | 0.988 | 0.387 | 0.434 | 0.683 | 0.910 | 0.921 | 0.987 |
HMSG | 0.604 | 0.498 | 0.843 | 0.889 | 0.887 | 0.990 | 0.345 | 0.265 | 0.563 | 0.929 | 0.925 | 0.988 |
MECCH | 0.825 | 0.819 | 0.946 | 0.936 | 0.933 | 0.993 | 0.527 | 0.529 | 0.746 | 0.953 | 0.951 | 0.994 |
MF2Vec | 0.904 | 0.907 | 0.986 | 0.958 | 0.956 | 0.990 | 0.667 | 0.670 | 0.861 | 0.963 | 0.962 | 0.995 |
ACM | |||||||||
Metapath | PAP | PSP | ALL | ||||||
Metric | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC |
MAGNN | 0.6064 | 0.6602 | 0.8754 | 0.7298 | 0.7649 | 0.8575 | 0.8095 | 0.8254 | 0.9480 |
HMSG | 0.6608 | 0.6024 | 0.8598 | 0.7656 | 0.7128 | 0.8302 | 0.8204 | 0.8160 | 0.9151 |
MECCH | 0.6683 | 0.6405 | 0.8431 | 0.7531 | 0.7067 | 0.8440 | 0.7980 | 0.7922 | 0.9194 |
MF2Vec | 0.8080 | 0.8013 | 0.9322 | 0.7631 | 0.7211 | 0.8436 | 0.8653 | 0.8610 | 0.9528 |
IMDB | |||||||||
Metapath | MAM | MDM | ALL | ||||||
Metric | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC |
MAGNN | 0.5241 | 0.5276 | 0.7072 | 0.5057 | 0.5180 | 0.6971 | 0.5685 | 0.5731 | 0.7493 |
HMSG | 0.4717 | 0.4587 | 0.6213 | 0.4226 | 0.3634 | 0.5554 | 0.4767 | 0.4357 | 0.6306 |
MECCH | 0.4149 | 0.3169 | 0.5887 | 0.4029 | 0.3601 | 0.5603 | 0.4639 | 0.4404 | 0.6449 |
MF2Vec | 0.5875 | 0.5733 | 0.7180 | 0.5851 | 0.5565 | 0.7142 | 0.6211 | 0.6117 | 0.7709 |
Model | Yelp | Movielens | DBLP | ACM | IMDB | Freebase | ||||||||||||
AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | AUC | Mi-F1 | Ma-F1 | |
GCN (Ours) | 0.8947 | 0.8200 | 0.8261 | 0.9074 | 0.8283 | 0.8366 | 0.9958 | 0.9631 | 0.9623 | 0.9528 | 0.8653 | 0.8610 | 0.7709 | 0.6211 | 0.6117 | 0.8701 | 0.7429 | 0.6991 |
Edge GAT | 0.8624 | 0.8041 | 0.7941 | 0.8629 | 0.8086 | 0.8180 | 0.9900.002 | 0.9280.007 | 0.9210.007 | 0.9500.005 | 0.8540.008 | 0.8490.010 | 0.7450.013 | 0.5790.023 | 0.5580.025 | 0.7950.010 | 0.6810.010 | 0.6140.026 |
GraphSAGE | 0.8296 | 0.7701 | 0.7771 | 0.8028 | 0.7506 | 0.7719 | 0.9830.005 | 0.9020.020 | 0.8960.022 | 0.9380.006 | 0.8160.009 | 0.8090.009 | 0.7020.018 | 0.5310.029 | 0.5010.031 | 0.7340.018 | 0.6400.021 | 0.4710.016 |
We evaluate the performance of these models across three downstream tasks: node classification, link prediction, and node clustering.
D. Ablation Study
D.1. Effect of Temperature ()
D.2. Graph Method
Table 14 compares GCN, EdgeGAT, and GraphSAGE to evaluate the impact of facet features as edge weights. GCN outperforms across all datasets, achieving AUC and F1 scores above 0.99 on DBLP and ACM by effectively leveraging edge weights to capture local and global relationships. While EdgeGAT models relationship importance using attention, its performance is less consistent, and GraphSAGE shows lower results due to its limited capacity to represent multi-facet relationships despite its efficiency. These findings highlight the significance of edge weights in facet-based graph learning.
F. Algorithms for Random Walk and Sub-graph Generation
This section describes the algorithms used for random walk and sub-graph generation in heterogeneous networks. The random walk algorithm explores paths within the graph, while the sub-graph generation algorithm utilizes these paths to construct sub-graphs based on specific node types.
F.1. Random Walk Algorithm
The first algorithm 2 performs a random walk on a network. It generates a path of a specified length starting from a given node. The path is built by selecting a random neighbor at each step. If a neighbor cannot be found or if the end node of the path does not match the desired node type, the algorithm returns None.
F.2. Sub-graph Generation Using Random Walk Paths
The second algorithm 3 generates a sub-graph from the network using multiple random walks. For each node, several random walks are performed to collect paths. Unique node pairs are recorded to avoid duplicates. The sub-graph is constructed by adding nodes and edges based on the collected paths.

### v3

@@ 4.5. Visualization
To evaluate embedding quality, we visualize node representations from various HGNNs using t-SNE (Van der Maaten and Hinton, 2008). Figure 7 shows t-SNE plots for HAN (Wang et al., 2019), MAGNN (Fu et al., 2020), HeCo (Wang et al., 2021), HMSG (Guan et al., 2023), MECCH (Fu and King, 2024), and MF2Vec, with colors representing author classes. MF2Vec demonstrates superior clustering within classes and clear decision boundaries, outperforming the more scattered intra-class distributions and weaker inter-class separations of other HGNNs.
Scatter plot showing the learned embeddings of nodes in DBLP, colored by label, demonstrating clear clustering by class.
4.6. Sensitivity to Random Walk
4.6.1. Effect of Random Walk Length.
In Table 9, increasing the random walk length potentially introduces more noise into the sampled paths, but the performance of MF2Vec remains stable across all tested values. Changes in AUC are minor, demonstrating that MF2Vec is insensitive to longer or noisier random walks. Based on these results, a length of 5 was used for experiments.
Dataset | 3 | 5 | 7 | 9 | |
DBLP | 0.990 0.008 | 0.992 0.002 | 0.988 0.010 | 0.982 0.009 | 0.010 |
ACM | 0.939 0.010 | 0.954 0.008 | 0.953 0.007 | 0.951 0.011 | 0.015 |
IMDB | 0.748 0.025 | 0.768 0.033 | 0.772 0.021 | 0.779 0.014 | 0.031 |
4.6.2. Effect of Random Walk Quality.
We simulate sparse environments by randomly removing 0%, 10% and 20% of node edges, then evaluate AUC and its standard deviation. As shown in Figure 8, MF2Vec shows the least performance variation, maintaining robust results despite increased noise and missing data. This demonstrates that MF2Vec leverages shared facet information to generalize well in incomplete or noisy settings.
Line chart comparing robustness of MF2Vec and other models when 0, 10, or 20 percent of edges are randomly removed. MF2Vec shows consistently higher performance across settings.
4.7. Analysis of Facet Balance
To further analyze the balance of facet distributions across node types, we visualized a facet-level attention heatmap as shown in Figure 9. Each column represents a specific node and each row corresponds to a facet. Darker colors indicate facets that play a more significant role in the node’s connections. For example, the paper node “P1612” mainly focuses on Facet 1, whereas the author node “A3958” distributes attention relatively evenly across all facets. This analysis confirms that some nodes concentrate on specific facets while others distribute attention more uniformly, with no strong bias overall. This visually validates the importance of modeling shared facets in heterogeneous graphs.
A heatmap showing facet-level attention across nodes in the DBLP dataset. Columns correspond to sampled nodes of different types, rows correspond to facets, and darker cells represent more influential facets.
4.8. Case Study
A path in the IMDB graph with nodes annotated by attributes such as rating, country, and language. All nodes share the language English, showing semantic grouping by facet.
Most graph datasets represent nodes by unique IDs, making it difficult to interpret learned embeddings in terms of real-world attributes. To address this, we conducted a qualitative case study to investigate whether the learned facets actually correspond to meaningful attributes. Based on the IMDB text data provided in MAGNN, we used SBERT to generate node embeddings and separated them into five explicit attributes shared by IMDB node types: rating, genre, country, language, and keywords. In this setting, we analyzed which attributes were shared among the intermediate nodes along a given path from a start node to an end node. For example, along the path “Horrible Bosses 2 → Penelope Wilton → All the Pretty Horses → Kevin Peter Hall → The Sorcerer’s Apprentice,” all nodes shared the attributes country (USA, UK), rating (around 6), and language (English). Notably, Gumbel-Softmax identified “language (English)” as the dominant facet. This shows that the connections in this path are not just simple Movie–Actor–Movie–Actor–Movie links, but are semantically grouped by language, thus demonstrating that MF2Vec learns interpretable and meaningful facets rather than mere latent ID embeddings.
4.9. Ablation Study
4.9.1. Effect of Temperature ()
A line plot showing MF2Vec performance across six datasets at different temperature values in Gumbel-Softmax. Performance varies slightly but remains stable overall.
4.9.2. Effect of Train Ratio Settings.
@@ 4.9.1. Effect of Temperature ()
As shown in Figure 12, MF2Vec consistently achieves the highest performance across varying training set ratios, demonstrating its robustness and adaptability. While the performance of all models declines as the training ratio decreases, MF2Vec remains superior in most cases and retains competitiveness in the rest. At the lowest ratio of 0.2, it occasionally shows comparable performance to other models. Only models achieving scores above a certain threshold on each dataset are displayed to ensure clearer visualization. These results highlight MF2Vec’s effectiveness in handling sparse training data.
A line chart showing classification performance of MF2Vec compared to baselines when varying the training set ratio from 0.8 to 0.2. MF2Vec consistently outperforms baselines.
4.9.3. Effect of Number of Facets.
@@
A plot showing MF2Vec’s performance trend as the number of facets increases. Best performance occurs at five facets, with degradation beyond that point.
5. Related Work
5.1. Heterogeneous Graph Neural Network
5.1.1. Relation-based HGNNs
5.1.2. Metapath-based HGNNs
5.2. Multiple-Vector Network Embedding
@@ 6. Conclusion
We introduced MF2Vec, a new heterogeneous graph neural network that learns node representations using multi-facet paths, independent of node types. MF2Vec consistently outperforms existing methods in node classification, link prediction, and clustering across six benchmark datasets. It demonstrates high stability against meta-path variations and offers improved computational efficiency. Our ablation results confirm the effectiveness of learnable facet selection. In future work, we plan to automate facet and path selection to further improve adaptability and performance. While MF2Vec shows strong generalization and scalability, its current path selection is based on random walks, which may miss some informative paths. Addressing this limitation is an important direction for future research. MF2Vec offers a simple and scalable solution for heterogeneous graph learning.
Acknowledgements.
This research was supported by National Research Foundation of Korea (NRF-2022M3J6A1063021).GenAI Usage Disclosure
The authors used Generative AI tools (ChatGPT, OpenAI) only for minor grammar checking and assistance in generating table formatting. No part of the scientific content, analysis, or results was produced by GenAI.